# End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding

### v3

@@ Abstract
Recent advances in AI foundation models have significant potential for lightening the clinical workload by mimicking the comprehensive and multi-faceted approaches used by medical professionals. In the field of radiation oncology, the integration of multiple modalities holds great importance, so the opportunity of foundational model is abundant. Inspired by this, here we present RO-LMM, a multi-purpose, comprehensive large multimodal model (LMM) tailored for the field of radiation oncology. This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM. In particular, to perform consecutive clinical tasks without error accumulation, we present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the consistency of handling clean inputs. We further extend this concept to LMM-driven segmentation framework, leading to a novel Consistency Embedding Segmentation (CESEG) techniques. Experimental results including multi-centre validation confirm that our RO-LMM with CEFTune and CESEG results in promising performance for multiple clinical tasks with generalization capabilities.
Index Terms:
Large multimodal model, Breast cancer, Clinical report, Radiotherapy target volume, Segmentation
I Introduction
@@
figureRO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology. The model seamlessly covers various tasks such as clinical report summarization, radiation treatment plan suggestion, and 3D target volume segmentation.
@@
Recently, the emergence of a new generation of AI models inspired by the large language models (LLMs), known as foundation models, marks a significant departure from previous paradigms [1]. These models are characterized by their massive scale and versatility, which stems from self-supervised training on a diverse array of data. These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [2, 3, 4, 5, 6, 7].
@@
These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision-making. Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [8]. Specifically, in contrast to these unimodal AIs, generalist medical AIs that synergistically combine foundation models can encompass a holistic understanding of clinical workflows, which can receive a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [9, 10, 11, 1, 12]. By understanding diverse data types and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
@@
In the field of radiation oncology, which the main focus of this article, the integration of multiple modalities holds great importance so that it is one of the most important clinical fields to evaluate the potential of foundation models. Accordingly, here we introduce RO-LMM, a prototype large multimodal model (LMM) specifically designed to support the clinical workflow in radiation oncology. In particular, this work significantly extends our related prior work called LLMSeg[13], which focused on multimodal segmentation. More specifically, RO-LMM enlarges the LLMSeg’s scope by tackling a broader range of clinical tasks in radiation oncology: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes. Additionally, it is capable of (2) proposing appropriate treatment plans from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed treatment plans. This multifaceted functionality of RO-LLM demonstrates a significant advances toward the expertise of clinical professionals.
@@
In training LLM to perform these sequential tasks from treatment plan suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each tasks which can potential leads to significant end-to-end performance loss. Accordingly, another important contribution of this work is the employment and extension of a Noisy Embedding Fine-Tuning (NEFTune) [14] which involves injecting uniform noise into embeddings during the training for each targeted task. More specifically, to further enhance the model’s applicability, we develop a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs. Furthermore, by expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in novel Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG). These advancements prevent from error propagation from each subsequent task and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation. As for the proof of the concept study, our RO-LLM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging. Our contributions can be summarized as:
@@
  • To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiation treatment plan suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
  • Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms baseline methods.
II Related Works
II-A Instruction Fine-tuning in LLM
@@
II-B Stabilizing LLM Fine-tuning with Noise
To enhance the robustness of language models against noisy input, various strategies have been explored. Approaches such as SMART [28], FreeLB [29], and R3F [30] employ adversarial training methods, introducing small Gaussian perturbations in embedding dimensions to optimize the model’s performance against noise. Aparting from the adversarial training, LSNR [31] takes a different approach by directly optimizing the smoothness of each layer of the language model. More recently, NEFTune [14] improves LLM performance during fine-tuning by simply adding random noise into the embedding vectors during training. In the context of this study, we extend NEFTune by incorporating the concept of consistency regularization. This extension aims to further improve the robustness and generalization capabilities of language models when faced with noisy input.
Figure 1: Schematics of RO-LMM training for three different tasks. (a) RO-LMM-S for clinical note summarization. (b) RO-LMM-P++ for radiation treatment plan suggestion. (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
II-C Language-driven Image Segmentation
Recent research in the field of image segmentation has been emerged to incorporate linguistic ability, such like language-driven semantic segmentation [32], open-vocabulary segmentation [33, 34], referring segmentation [35], and reasoning segmentation [7]. Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable. For instance, LViT [36] and ConTEXTualNet [37] introduce text-driven chest X-ray radiography segmentation.
Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [38, 39, 40]. Recent research on clinical context-aware breast cancer radiotherapy delineation [13] has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce. By extending [13], we incorporate treatment plans generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
III Methods
@@
III-A Consistency Embedding Fine-tuning for Clinical LMM
To realize the multi-purpose LMM with expertise in clinical report summarization and radiation treatment plan suggestion, we conduct instruction fine-tuning for LLaMA2 [23]. Considering the nuanced differences in the intended objective of each task, we adopt separate training strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
Specifically, we train a summary expert using collected raw clinical report and summary notes. During inference, the summary expert receives raw clinical reports, just like in the training scenario. However, for the plan expert, there is a discrepancy between the training and inference scenarios. In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure I. However, at the inference phase, our model takes the generated notes from the trained summary expert. To deal with the input domain differences for the training and inference time, Noisy Embedding Fine-Tuning (NEFTune) [14] which inject uniform noise into embedding could be an effective naive solution to handle noisy inputs in this task. However, a crucial consideration arises from the nature of the generated notes, since some of them may lie closer to clean inputs (collected notes) and the others deviate towards noisy inputs. To address this, it is essential to train the model to handle both clean and noisy inputs. To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++111‘+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively. More details are as follows.
Consistency Embedding Fine-Tuning. To enhance the performance of instruction fine-tuning, NEFTune  [14] involves injecting a random noise vector into embeddings during the training process as follows:
@@ III-A Consistency Embedding Fine-tuning for Clinical LMM
where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ is perturbed embedding, $f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample. They demonstrated that the effectiveness of incorporating a random noise vector adjusted based on token length, which yields robust results in most fine-tuning scenarios.
Figure 2: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module. These module outputs are aligned through (c) multimodal alignment module.
To bolster the model’s capability in handling both noisy and clean inputs for treatment plan suggestion task, we incorporate a regularization loss to encourage consistency as follows:
@@
where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$. The metric function $d$ can be any capable of measuring distance in a specific space. The function $\mathcal{T}$ performs detokenization to generate sentences.
By introducing the consistency in the detokenized domain through $\mathcal{T}$, our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs. Since solely minimizing distance in the same embedding space may weaken the robustness and pose limitations on enforcing semantic similarity, to capture textual similarity between sentences, another contribution of our work is to calculate the distance in the feature space of SentenceBERT [41]. i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using pretrained model such as SentenceBERT [41].
III-B LMM-assisted 3D Target Volume Segmentation
For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LLM for context-aware 3D segmentation introduced in LLMSeg[13]. LLMSeg demonstrated significant performance improvement in multimodal radiotherapy target segmentation compared to unimodal models. Additionally, it showed the ability to align target contours with provided clinical text information. However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated treatment plan from the previous summarization step. To address this noise, we introduce two novel modules: Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG). These modules enhance the target volume segmentation network by handling noise and promoting consistency. The details are as follows.
Noise Embedding Segmentation (NESEG). Firstly, we introduce the noise augmentation technique, dubbed as NESEG, for the target volume segmentation task. We refer to the corresponding RO-LLM as RO-LMM-SEG+. Specifically, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, leading to the follow loss function:
@@ III-B LMM-assisted 3D Target Volume Segmentation
Consistency Embedding Segmentation (CESEG). We further adapt the consistency regularization module, dubbed as CESEG, for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan. The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 2.
Apart from the original concept of CEFTune, CESEG modifies the consistency regularization module for the multimodal segmentation task, by combining CEFTune and the text prompt tuning. Once the text prompt-prepended noisy embeddings of treatment plan ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. Specifically, the loss function for CESEG is formulated as:
@@
We employ the 3D Residual U-Net [42] as for the backbone structure within the image module, as shown in (a). In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the treatment plan embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input. In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [43]). For realizing multimodal alignment within the 3D segmentation framework, we adapt the idea of light-weight text prompt tuning to transfer the linguistic capability of LLM, by following the concepts introduced in [13, 44, 45]. Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module.
IV Experiments
IV-A Dataset
To train our model, we collected internal data cohort composed of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center. For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 599 patients with their 3D CT scans and corresponding ground truth masks. To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital. The detail of dataset is described in Table I and Supplementary Material. We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals. This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
IV-B Implementation details
For instruction fine-tuning for RO-LMM-S and RO-LMM-P, we use the LLaMA-2-7B-Chat [23] model as a baseline. The maximum context length is set at 4096, and the batch size is set to 2. To train RO-LMM-SEG, we employ the 3D Residual U-Net [42] using the open-source library MONAI [^2] https://monai.io/. For incorporating the textual information within the multimodel segmentation model, we use the LLaMA-2-7B-Chat model initialized with the pre-trained checkpoint of RO-LMM-P++. We set the batch size as 1. We use the AdamW [46] optimizer for all the tasks, with a learning rate of 5e-5 until reaching 3 epochs for both RO-LMM-S and RO-LMM-P, and 1e-4 until reaching 100 epochs for RO-LMM-SEG, leveraging 4 NVIDIA A6000 GPUs for each task. The hyper-parameter $\lambda$ for CEFTune is set to 1 in all of tasks, and for consistency regularization we adopt the variants of SBERT [41] which is trained on PubMed [^3] https://www.ncbi.nlm.nih.gov/pubmed/ dataset, called PubMedBERT. [^4] https://huggingface.co/NeuML/pubmedbert-base-embeddings.
TABLE I: Training data details. CRS: Clinical Report Summarization. TPS: Treatment Plan Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
@@ IV-B Implementation details
CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 60 81
TPS Clinical Note Treatment Plan
PTS Treatment Plan &\&& CT Scan Segmentation 593 79 81
TABLE II: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
Model Method Internal Validation (N=60) External Validation (N=81)

### v4

@@ Abstract
Recent advances in AI foundation models have significant potential for lightening the clinical workload by mimicking the comprehensive and multi-faceted approaches used by medical professionals. In the field of radiation oncology, the integration of multiple modalities holds great importance, so the opportunity of foundational model is abundant. Inspired by this, here we present RO-LMM, a multi-purpose, comprehensive large multimodal model (LMM) tailored for the field of radiation oncology. This model effectively manages a series of tasks within the clinical workflow, including clinical context summarization, radiotherapy strategy suggestion, and plan-guided target volume segmentation by leveraging the capabilities of LMM. In particular, to perform consecutive clinical tasks without error accumulation, we present a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which boosts LMM’s robustness to noisy inputs while preserving the consistency of handling clean inputs. We further extend this concept to LMM-driven segmentation framework, leading to a novel Consistency Embedding Segmentation (CESEG) techniques. Experimental results including multi-centre validation confirm that our RO-LMM with CEFTune and CESEG results in promising performance for multiple clinical tasks with generalization capabilities.
keywords:
Large multimodal model , Breast cancer , Clinical report , Radiotherapy target volume , Multimodal segmentation
journal: Medical Image Analysis
@@
figureRO-LMM as an assistant large multimodal model (LMM) in the field of radiation oncology. The model seamlessly covers various tasks such as clinical report summarization, radiation radiotherapy strategy suggestion, and 3D target volume segmentation.
1 Introduction
@@
Recently, the emergence of a new generation of AI models inspired by large language models (LLM), known as foundation models, marks a significant departure from previous paradigms [36]. These models are characterized by their massive scale and versatility, which stems from self-supervised training on a diverse array of data. These foundation models are now capable of achieving state-of-the-art performance (SOTA) in a wide range of domains, including tasks such as multimodal reasoning, image-to-text generation, image captioning, and text-guided image segmentation [3, 11, 13, 28, 33, 26].
@@
These characteristics signify a potential paradigm shift in how AI can be integrated into medical practices, which inherently rely on multimodal information for comprehensive clinical decision making. Furthermore, this gives an opportunity of overcoming limitations of now over 500 FDA-approved AI models, which are mainly specialized for a specific task with unimodal information [22]. Specifically, in contrast to these unimodal AIs, generalist medical AIs that synergistically combine foundation models can encompass a holistic understanding of clinical workflows, which can receive a variety of medical data, including imaging modalities, electronic health records, laboratory results, genomics, and even clinical reports [43, 41, 51, 36, 47]. By understanding various types of data and their interrelationships, multimodal AIs would facilitate more accurate diagnoses, personalized treatment development, and a reduction in medical errors by providing a comprehensive view of patient data.
@@
In radiation oncology, the primary focus of this article, the integration of multiple modalities is crucial, making it one of the most important clinical fields for evaluating the potential of foundation models. Accordingly, here we introduce RO-LMM, a prototype large multimodal model (LMM) specifically designed to support the clinical workflow in radiation oncology. In particular, this work significantly extends our prior related work called LLMSeg[37], which focused on multimodal segmentation. More specifically, RO-LMM enlarges the LLMSeg’s scope by tackling a broader range of clinical tasks in radiation oncology: (1) It efficiently summarizes extensive patient histories and examination results into concise but informative clinical notes. Additionally, it is capable of (2) proposing appropriate radiotherapy strategies from a clinical expert perspective and (3) delineating radiation target volumes on 3-dimensional (3D) computed tomography (CT) scans consistent with the proposed radiotherapy strategies. This multifaceted functionality of RO-LMM demonstrates significant advances in supporting the expertise of clinical professionals.
@@
In training LLM to perform these sequential tasks from radiotherapy strategy suggestion to target volume segmentation, we found that there exist potentials for error accumulation from each task which can potential leads to significant end-to-end performance loss. Accordingly, another important contribution of this work is the employment and extension of a Noisy Embedding Fine-Tuning (NEFTune) [20] which involves injecting uniform noise into embeddings during the training for each targeted task. More specifically, to further enhance the model’s applicability, we develop a novel Consistency Embedding Fine-Tuning (CEFTune) technique, which add regularization loss to enforce the consistency between the prediction given noisy and clean inputs. Furthermore, by expanding beyond text-related tasks, we apply these concepts to 3D segmentation tasks, resulting in novel Noisy Embedding Segmentation (NESEG) and Consistency Embedding Segmentation (CESEG). These advancements prevent error propagation between subsequent tasks and collectively contribute to a marked improvement in the end-to-end model’s generalization capabilities in both internal and external validation.
As for the proof of the concept study, our RO-LMM framework was adopted for breast cancer study, one of the high-prevalence cancer types requiring relatively standardized radiotherapy based solely on CT imaging. Our contributions can be summarized as:
  • 1.
@@
  • 2.
    To prevent from potential error accumulation throughout consecutive clinical tasks such as clinical context summarization, radiotherapy strategy suggestion, and plan-guided target volume segmentation, we explore noise augmentation and consistency, and propose a novel training approach, such as CEFTune, NESEG, and CESEG that significantly enhance the robustness of our method.
  • 3.
    Through experiments on various validation settings with real clinical data of breast cancer patients, we demonstrated that RO-LMM outperforms Default methods.
Fig. 1: Schematics of RO-LMM training for three different tasks. (a) RO-LMM-S for clinical note summarization. (b) RO-LMM-P++ for radiotherapy strategy suggestion. (c) RO-LMM-SEG++ for plan-guided target volume segmentation.
2 Related Works
2.1 Instruction Fine-tuning in LLM
@@
2.2 Stabilizing LLM Fine-tuning with Noise
To improve the robustness of language models against noisy input, various strategies have been explored. Approaches such as SMART [21], FreeLB [54], and R3F [1], LSNR [18] utilize adversarial training by introducing minor Gaussian perturbations in embedding dimensions, optimizing model performance in noisy environments. More recently, NEFTune [20] improves LLM performance during fine-tuning by simply adding random noise into the embedding vectors during training. In the context of this study, we extend NEFTune by incorporating the concept of consistency regularization. This extension aims to further improve the robustness and generalization capabilities of language models when faced with noisy input.
2.3 Language-driven Image Segmentation
Recent research in the field of image segmentation has emerged to incorporate linguistic ability, such like language-driven semantic segmentation [27], open-vocabulary segmentation [12, 31], referring segmentation [49], and reasoning segmentation [26]. Language-driven segmentation has triggered a paradigm shift in medical domain, where multimodal knowledge is inevitable. For instance, LViT [30] and ConTEXTualNet [19] introduce text-driven chest X-ray radiography segmentation. Target volume segmentation in the field of radiation oncology is more challenging, due to its intrinsic need for considerations of the clinical aspects beyond the image, such as overall cancer stage, treatment aim, pathological findings, and so on [6, 7, 17]. Recent research on clinical context-aware breast cancer radiotherapy delineation [37] has demonstrated that the multimodal AI outperforms the traditional unimodal AI by a substantial margin, particularly when labeled datasets are scarce. By extending [37], we incorporate radiotherapy strategies generated by LMM directly from clinical reports, which aligns seamlessly with actual clinical workflows in radiation oncology.
3 Methods
@@
3.1 Consistency Embedding Fine-tuning for Clinical LMM
To realize the multi-purpose LMM with expertise in clinical report summarization and radiotherapy strategy suggestion, we conduct instruction fine-tuning for LLaMA2 [46]. Considering the nuanced differences in the intended objective of each task, we adopt separate training strategies to acquire task-specific expertise, namely RO-LMM-S (summary expert) and RO-LMM-P (plan expert).
Specifically, we train a summary expert using collected raw clinical report and summary notes. During inference, the summary expert receives raw clinical reports, just like in the training scenario. However, for the plan expert, there is a discrepancy between the training and inference scenarios. In other words, we use the training set made up of collected summary notes instead of generated notes from the summary expert, mainly due to cost concerns and the inherent nature of our framework, which generates output sequentially as illustrated in Figure End-to-End Breast Cancer Radiotherapy Planning via LMMs with Consistency Embedding. However, at the inference phase, our model takes the generated notes from the trained summary expert. To deal with the input domain differences for the training and inference time, Noisy Embedding Fine-Tuning (NEFTune) [20] which inject uniform noise into embedding could be an effective naive solution to handle noisy inputs in this task. However, a crucial consideration arises from the nature of the generated notes, since some of them may lie closer to clean inputs (collected notes) and the others deviate towards noisy inputs. To address this, it is essential to train the model to handle both clean and noisy inputs. To preserve the robustness facilitated by NEFTune while enforcing consistency between the prediction given clean and noisy inputs, we introduce Consistency Embedding Fine-Tuning (CEFTune), resulting in RO-LMM-P++111‘+’ , ‘++’ denotes the adoption of NEFTune, and CEFTune, respectively. More details are as follows.
3.2 Consistency Embedding Fine-Tuning.
To enhance the performance of instruction fine-tuning, NEFTune  [20] involves injecting a random noise vector into embeddings during the training process as follows:
@@ 3.1 Consistency Embedding Fine-tuning for Clinical LMM
where ${\mathbf{x}^{\text{emb}}_{\text{txt}}}\in\mathbb{R}^{B\times L\times C}$ is the embedding of data, $B$ denotes batch size, $L$ is token length, $C$ is embedding dimension, $\alpha$ is a tunable parameter. The perturbed embedding is denoted as $\tilde{\mathbf{x}}$, and $\epsilon$ represents a random noise vector, drawn from a uniform distribution $\mathcal{U}(-1,1)$. Additionally, $\mathbb{E}$ represents the expectation. $f_{\theta}(\cdot)$ represents the model parameterized by $\theta$, ${\mathbf{y}_{\text{txt}}}$ is the label of text sample. They demonstrated that the effectiveness of incorporating a random noise vector adjusted based on token length, which yields robust results in most fine-tuning scenarios.
To bolster the model’s capability in handling both noisy and clean inputs for radiotherapy strategy suggestion task, we incorporate a regularization loss to encourage consistency as follows:
@@
where $\theta^{-}$ represents the model with stopped gradient, and $d(\cdot,\cdot)$ is used to quantify the discrepancy between $f_{\theta}({\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}})$ and $f_{\theta^{-}}({\mathbf{x}^{\text{emb}}_{\text{txt}}})$. The metric function $d$ can be any capable of measuring distance in a specific space. The function $\mathcal{T}$ performs detokenization to generate sentences. By introducing the consistency in the detokenized domain through $\mathcal{T}$, our objective is to preserve the robustness property of NEFTune while introducing semantic similarity through the integration of consistency between the output of noisy inputs and clean inputs. Since solely minimizing distance in the same embedding space may weaken the robustness and pose limitations on enforcing semantic similarity, to capture textual similarity between sentences, another contribution of our work is to calculate the distance in the feature space of SentenceBERT [42]. i.e., $d(\mathcal{T}(x),\mathcal{T}(y))=1-s(\mathcal{T}(x))^{\top}s(\mathcal{T}(y))/||s(\mathcal{T}(x))||_{1}||s(\mathcal{T}(y))||_{1}$, where $s(\cdot)$ is the projection by using a pretrained model such as SentenceBERT [42].
Fig. 2: Schematics of RO-LMM-SEG++ for plan-guided 3D target volume segmentation task, which composed of (a) image module and (b) text module. These module outputs are aligned through (c) multimodal alignment module.
3.3 LMM-assisted 3D Target Volume Segmentation
For incorporating the textual information into target volume segmentation framework, we expand the concept of prompt tuning of LLM for context-aware 3D segmentation introduced in LLMSeg[37]. LLMSeg demonstrated significant performance improvement in multimodal radiotherapy target segmentation compared to unimodal models. Additionally, it showed the ability to align target contours with provided clinical text information. However, in the extended clinical workflow proposed here, the LMM-assisted segmentation model receives noisy text descriptions of the generated radiotherapy strategy from the previous summarization step. To address this noise, we introduce two novel modules: Noise Embedding Segmentation (NESEG) and Regularized Consistency Embedding Segmentation (CESEG). These modules enhance the target volume segmentation network by handling noise and promoting consistency. The details are as follows.
3.4 Noise Embedding Segmentation (NESEG).
Firstly, we introduce the noise augmentation technique, dubbed as NESEG, for the target volume segmentation task. We refer to the corresponding RO-LMM as RO-LMM-SEG+. Specifically, by extending aforementioned idea of NEFTune, we inject a random noise into input data embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ to improve the segmentation model robustness for the generated noisy plan, leading to the follow loss function:
@@ 3.3 LMM-assisted 3D Target Volume Segmentation
3.5 Consistency Embedding Segmentation (CESEG).
We further adapt the consistency regularization module, dubbed as CESEG, for generalizing our segmentation model to both the generated noisy plan and the clean ground truth plan. The schematic of our multimodal model utilizing CESEG, denoted as RO-LMM-SEG++, is illustrated in Figure 2.
Apart from the original concept of CEFTune, CESEG modifies the consistency regularization module for the multimodal segmentation task, by combining CEFTune and the text prompt tuning. Once the text prompt-prepended noisy embeddings of radiotherapy strategy ${\mathbf{\tilde{x}}^{\text{emb}}_{\text{txt}}}$ are inputted to the multimodal model, the output embedding is also regularized with respect to the clean embeddings ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$. Specifically, the loss function for CESEG is formulated as:
@@
We employ the 3D Residual U-Net [8] as for the backbone structure within the image module, as shown in (a). In the text module, as shown in (b), we employ pre-trained RO-LMM-P++ as for the LLM module $f_{\theta^{*}}$ for producing text embedding ${\mathbf{y}^{\text{emb}}_{\text{txt}}}$ given the radiotherapy strategy embedding ${\mathbf{x}^{\text{emb}}_{\text{txt}}}$ as input. In the multimodal alignment module, as illustrated as (c), we adopt both self-attention and cross-attention mechanisms in an interactive manner (image-to-text and text-to-image) for each $l$-th skip-connection layer, by following the concept of promptable segmentation borrowed from Segment Anything Model (SAM) [24]). For realizing multimodal alignment within the 3D segmentation framework, we adapt the idea of light-weight text prompt tuning to transfer the linguistic capability of LLM, by following the concepts introduced in [37, 23, 53]. Therefore, entire parameters of the LLM module $f_{\theta^{*}}$ are frozen, while the learnable text prompts $\mathbf{z}_{\phi}$ and $l$-th linear layers are exclusively optimized within the text module.
Table 1: Training data details. CRS: Clinical Report Summarization. RSS: Radiotherapy Strategy Suggestion. PTS: Plan-guided Target Segmentation. US: Ultrasound. Path: Pathology.
@@
CRS MRI &\&& US &\&& Path Reports Clinical Note 5,674 121 81
RSS Clinical Note Radiotherapy Strategy
PTS Radiotherapy Strategy &\&& CT Scan Segmentation 593 79 81
4 Experimental settings
4.1 Dataset
To train our model, we collected an internal dataset consisting of 5,674 breast cancer patients treated at the Department of Radiation Oncology at Yonsei Cancer Center. For training the plan-guided 3D target volume segmentation network, we utilized multimodal data from 593 patients, including their 3D CT scans and corresponding ground truth masks. For the text-related dataset, we established a validation set with 121 patients. Due to the limited availability of image data compared to text data, the validation size for the segmentation task was set to 79 patients.
To evaluate the model performance on cross-centre datasets, we further acquired external data cohort composed of 81 patients treated at the Department of Radiation Oncology at Yongin Severance Hospital. The detail of dataset is described in Table 1 and A. We included patients who received their initial diagnosis of breast cancer and subsequently underwent radiation therapy following curative surgery, while excluding individuals with recurrent or metastatic breast cancer in both hospitals. This study was approved by the Institutional Review Board (IRB) of each participating hospitals.
Table 2: Quantitative comparison for clinical note summarization. Vanilla: the instruction fine tuning. CI: confidence interval.
Model Method Internal Validation (N=121) External Validation (N=81)